OpenStudy (anonymous):

Can someone help me with this equation for A : 2A/3 = 8 + 4A it's for nursing exam

OpenStudy (anonymous):

don't tell answer please explain.

OpenStudy (mathmale):

A warm welcome to OpenStudy! I assume you want to solve this equation for the variable A. Is that correct? Next time, please share the instructions that came with this problem. To solve this equation for A, we need to consolidate all A terms on the left side and all terms without A on the right side. 2A/3 = 8 + 4A includes a fraction (2/3) in the first term. It would be easier to solve this equation were we to eliminate that fraction. To do so, multiply all 3 terms of this equation by 3. Would you do that now, please? Type your result here.

OpenStudy (mathmale):

Hints: 3(8) = 24; 3(4A) = 12A; \[3\frac{ 2A }{ 3 }=~?\]

OpenStudy (anonymous):

2a

OpenStudy (anonymous):

gosh your good im still confused

OpenStudy (anonymous):

okay I get it alittle

OpenStudy (mathmale):

Ask yourself: What are we trying to do here, by multiplying everything by 3?

OpenStudy (mathmale):

\[3(\frac{ 2A }{ 3 })=8+4A)~becomes~2A=24+12A\]

OpenStudy (mathmale):

What next? Always keep our goal in mind: We want to find a value for A that satisfies the original equation.

OpenStudy (anonymous):

-2.4

OpenStudy (mathmale):

That looks correct. How would you test your answer, to ensure that it really is a solution of the original equation?

OpenStudy (anonymous):

replace the answer at the A point

OpenStudy (mathmale):

Yes: where you have A in the original equation, replace it with (-2.4). Be certain to use parentheses as I've done here. Would you try checking your answer now, by substitution?

OpenStudy (mathmale):

\[\frac{ 2A }{ 3 }=8+4A\] becomes \[\frac{ 2(-2.4) }{3 }=8+4(-2.4)\]

OpenStudy (mathmale):

I'm leaving it up to you to determine whether or not this equation is true. if it is true, then our A=-2.4 is the solution of the original equation.

OpenStudy (mathmale):

I'd appreciate a response from you.

OpenStudy (anonymous):

Yes I hear ya Ithis thing is slow to respond it is true.